### v1

# High-Fidelity SLAM Using Gaussian Splatting with Rendering-Guided Densification and Regularized Optimization

Abstract
We propose a dense RGBD SLAM system based on 3D Gaussian Splatting that provides metrically accurate pose tracking and visually realistic reconstruction. To this end, we first propose a Gaussian densification strategy based on the rendering loss to map unobserved areas and refine reobserved areas. Second, we introduce extra regularization parameters to alleviate the “forgetting” problem in the continuous mapping problem, where parameters tend to overfit the latest frame and result in decreasing rendering quality for previous frames. Both mapping and tracking are performed with Gaussian parameters by minimizing re-rendering loss in a differentiable way. Compared to recent neural and concurrently developed gaussian splatting RGBD SLAM baselines, our method achieves state-of-the-art results on the synthetic dataset Replica and competitive results on the real-world dataset TUM.
Figure 1: Overview of the method. Our method takes RGBD frames as inputs. During mapping, when given a posed RGBD frame, we first render the opacity image, color image and depth image. Then we compare them with the ground truth to densify the existed map. During tracking, we minimize the color and depth re-rendering loss to optimize the camera pose.
I Introduction
Dense visual SLAM with RGBD inputs is essential for many downstream tasks in mobile robots, AR/VR and robot manipulation. Traditional SLAM systems have focused mainly on camera tracking [1] and geometric surface reconstruction [2]. However, surface appearance reconstruction, which contains rich information for scene understanding, is often lacking in traditional SLAM systems. Motivated by the success of NeRF [3] on novel view synthesis, recent RGBD SLAM methods use volumetric structures [4, 5, 6, 7] or points [8] as the map representation. Combined with neural networks as decoders, these methods can achieve both texture and geometry reconstruction. However, limited by the expensive sampling along the ray in the optimization of NeRF, the above mentioned methods cannot render full-resolution images with high quality (commonly only 0.5% pixels are used for mapping [5], leading to low quality image rendering; see fig. 3 and table I).
The recent work 3D Gaussian Splatting (3DGS) [9] provides a more efficient way for novel view synthesis. Instead of expensive sampling along the ray, 3DGS relies on rasterization for rendering, which accelerates the mapping process so as to handle full-resolution images. The original 3DGS needs prior camera poses, which are often estimated by Structure-from-Motion [10]. In this work, we extend 3DGS in an SLAM framework, removing the need for prior camera poses. Our method processes sequential RGBD frame inputs, simultaneously optimizing Gaussian parameters and estimating camera poses. Though there are some concurrent works combining 3DGS with SLAM [11, 12], we show our method achieves better or competitive results compared to recent baselines.
Our first contribution is the introduction of a novel densification strategy based on rendering. We directly densify the map utilizing rendering loss, enabling us to effectively map unobserved areas and enhance the rendering quality of reobserved regions.
Our second contribution is the regularized optimization during the mapping. The continuous mapping suffers from the “forgetting” problem, which means that Gaussian parameters tend to overfit the latest frame and lead to decrease of reconstruction quality for previous frames (cf. Figure 2). To alleviate the forgetting problem, we introduce extra parameters to supervise the learning process, and the experiment shows that our method can preserve the rendering quality of previously visited areas.
Experiments show that our method achieves state-of-the-art reconstruction quality and tracking accuracy on Replica dataset and competitive results on the real-world dataset TUM-RGBD. In the ablation study, we show the effectiveness of our method. We also find that continuous mapping can reconstruct high-quality images even if there exists motion blur, which may inspire the 3DGS reconstruction with blurry inputs.
II Related Work
II-A Visual SLAM
Conventional visual SLAM methods are divided into direct and indirect methods. Indirect methods (e.g., ORB-SLAM [1]) first detect points of interest and attach them by feature descriptors. The tracking is conducted by feature matching and minimizing the re-projection error across frames. Direct methods (e.g., DSO [13]) consume raw pixels and construct photometric loss for tracking. These conventional visual SLAM methods focus more on tracking instead of the scene reconstruction. Since NeRF [14] shows powerful scene reconstruction abilities, recent work applies neural representation in the SLAM to produce more complete and photorealistic scene reconstruction. For example, iMAP [7] uses a neural network to represent the scene; NICE-SLAM [4] and later work [5, 6] combine neural networks and voxel grids to improve reconstruction quality. Thanks to the generalization capabilities of neural networks, these neural representations have more powerful abilities to fill holes and construct more complete scenes compared to explicit representations. Although neural representations achieve good results in geometric scene reconstruction, their performance in recovering high-fidelity scenes is not satisfactory enough. In this work, we use parameterized Gaussians [9] as the map representation and achieve a better photorealistic reconstruction. Concurrent with our work, some SLAM systems are also based on Gaussian Splatting: MonoGS [15] expands the map by random sampling; Gaussian-SLAM [16] divides the scene into many submaps; Photo-SLAM [17] conducts tracking by minimizing the reprojection error; SplaTAM [11] and GS-SLAM [12] densify the map only considering the unobserved regions. Our method expands the map by considering filling holes in unobserved areas and refining reobserved areas, which can effectively improve the reconstruction quality. In addition, our regularized optimization can effectively alleviate the forgetting problem during mapping and preserve details of previously visited areas.
II-B Photo-realistic reconstruction
NeRF [14] opens a convenient way for photo-realistic reconstruction from only 2D images. Based on neural networks and volumetric differentiable rendering, NeRF can render novel photorealistic views that were not observed in the inputs. Many follow-up works promote original NeRF by accelerating training speed [18], improving rendering quality [19], etc. Some works [20] also reveal that neural networks are not necessary for novel view synthesis. As opposed to Gaussian Splatting [9], however, all the above-mentioned methods adopt expensive ray-marching when calculating one pixel, resulting in low rendering speed (lower than 15 fps). Gaussian Splatting uses parameterized Gaussians (as introduced in section III) as rendering primitives. Instead of expensive ray-marching, Gaussian Splatting renders the image by tile-based rasterization, which can achieve much faster rendering speed (around 100 fps). The original Gaussian Splatting needs known camera poses (commonly achieved by SfM [10]) and optimizes Gaussian parameters offline. In this work, we consider sequential RGBD frame inputs and optimize parameters online, which releases the requirement of known camera poses and brings the possibility of online applications such as exploration.
III Preliminary: Gaussian Splatting Rendering
We use 3D Gaussians as the map representation primitives for the color and depth image rendering. As described in [9], each Gaussian consists of parameters
$\mathbf{\Theta}_{i}\mathrel{\mathop{:}}=\{\bm{\mu}_{i},\bm{s}_{i},\bm{r}_{i},\bm{c}_{i},o_{i}\},$ (1)
where Gaussian means $\bm{\mu}_{i}\in\mathbb{R}^{3}$, scales $\bm{s_{i}}\in\mathbb{R}^{3}$, rotation quaternion $\bm{r}_{i}\in\mathbb{R}^{4}$, RGB color $\bm{c}_{i}\in\mathbb{R}^{3}$, and opacity $o_{i}\in\mathbb{R}$. All of these parameters are optimizable during training.
When rendering the image at some camera pose $T$, the 3D Gaussians in the field of view are projected onto the image plane as 2D Gaussians with mean $\bm{\mu}_{i}^{\prime}$ and covariance $\bm{\Sigma}_{i}^{\prime}$. To render the color $C$ and depth $D$ of the selected pixel $x$ on the image plane, we blend $N$ ordered (by depth) 2D Gaussians overlapping the pixel:
$C=\sum_{i\in N}\bm{c}_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j})$ (2)
$D=\sum_{i\in N}z_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j})$ (3)
with
$\alpha_{i}=o_{i}\cdot\exp{[-\frac{1}{2}(\bm{x}-\bm{\mu^{\prime}})(\bm{\Sigma}^{\prime})^{-1}(\bm{x}-\bm{\mu^{\prime}})]}.$ (4)
Though eq. 2 is the same as the volume rendering used in the classical NeRF [14], rasterization-based rendering is much more efficient. Because rasterization allows us to project and render objects in the space directly instead of expensive random sampling along the ray used in a full volumetric representation.
IV Methods
Figure 1 shows an overview of our method. Given RGBD frames and estimated camera poses, we update the map by comparing the rendered images and the ground truth to identify unobserved regions and areas requiring refinement. Regularization terms are incorporated into the optimization process to mitigate the issue of forgetting during mapping (section IV-A). We track the camera pose in the Gaussian map by minimizing color and depth re-rendering loss (section IV-B).
IV-A Mapping
Gaussian Densification. Different from initializing the Gaussian parameters with all images in the original 3D Gaussian Splatting [9], we process sequential RGBD frame inputs. For the latest RGBD input frame with an estimated pose (section IV-B), we need to add new Gaussians to the map, either to complete unobserved regions or to improve the rendering quality of previously mapped regions.
First, for the unobserved regions, we need to add new Gaussians to fill holes. We judge holes by rendering opacity. That is, when receiving a new frame, we render the opacity image at the estimated pose by
$O=\sum_{i\in N}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j})$ (5)
In the opacity image, if the rendered value is below some predefined opacity threshold $\tau_{opa}$ (i.e., $O(x,y)<\tau_{\mathrm{opa}}$), we need to densify the corresponding regions to fill holes.
Second, due to e.g. occlusion and illumination changes, the rendered image is often view-dependent; in addition, the optimized Gaussian parameters may get trapped into local minima to fit certain frames. To account for such view-dependent errors, we propose a kind of rendered error guided densification. When giving the estimated pose, we render the color image $\hat{C}$ and the depth image $\hat{D}$ at this pose; the current reference color image and the depth image are $C$ and $D$. We add new Gaussians when the error between the rendered image and the input image is large, where the color error and depth error are $E_{\mathrm{color}}(x,y)=|\hat{C}(x,y)-C(x,y)|$ and $E_{\mathrm{depth}}(x,y)=|\hat{D}(x,y)-D(x,y)|$ respectively. We will densify those regions where $E_{\mathrm{color}}(x,y)>\tau_{\mathrm{color}}$ or $\frac{E_{\mathrm{depth}}(x,y)}{D(x,y)}>\tau_{\mathrm{depth}}$, where $\tau_{\mathrm{color}}$ and $\tau_{\mathrm{depth}}$ are predefined thresholds.
In summary, we densify the regions where $\big{(}O(x,y)<\tau_{\mathrm{opa}}\big{)}\ |\ \big{(}E_{\mathrm{color}}(x,y)>\tau_{\mathrm{color}}\big{)}\ |\ \big{(}\frac{E_{\mathrm{depth}}(x,y)}{D(x,y)}>\tau_{\mathrm{depth}}\big{)}$. With depth images, we project corresponding depths to the space directly. The experiment (table IV) shows our densification strategy can significantly improve the map quality.
Continuous Mapping. We optimize Gaussian parameters by minimizing the loss $\ell_{1}$ between the rendered color and depth image and the input frames; in addition, following the original implementation in [9], we add an extra $\mathrm{SSIM}$ term.
$\mathcal{L}_{\mathrm{mapping}}=$ $\lambda_{\mathrm{color}}|\hat{C}-C|+\lambda_{\mathrm{depth}}|\hat{D}-D|$
$+\lambda_{\mathrm{SSIM}}\text{SSIM}(\hat{C},C)$ (6)
As a continuous learning problem, the incremental mapping in our work also suffers from the forgetting problem [21]. That is, the same Gaussian might be “seen” in multiple camera poses, but the Gaussian parameters tend to be optimized (overfitted) to fit the latest frame, resulting in a performance drop for previous frames (as demonstrated in fig. 2).
Figure 2: Illustration of the forgetting problem in the context of continual mapping based on Gaussians. The Gaussians colored by yellow are shared by camera0 and camera1. However, these Gaussians tend to be optimized to overfit the latest frame camera1, resulting in drop of reconstruction quality for previous frames.
To alleviate the overfitting problem, prior work [4, 5, 6] maintains a keyframe buffer. When optimizing the latest frame, some previous keyframes are selected from the buffer based on some criteria such as overlap ratio with the current frame. The selected keyframes will then be optimized together with the current frame. In the original 3D Gaussian Splatting, all frames are queued and picked randomly for many iterations to optimize Gaussian parameters.
We can, of course, pick many previous frames whenever we optimize the current frame; however, optimizing with multiple frames will lead to more time consumption, which is not suitable for online continual mapping. To prevent Gaussian parameters from forgetting, we add an extra regularization term $\mathcal{L}_{\mathrm{reg}}$ when mapping:
$\mathcal{L}_{\text{reg}}=\sum_{i\in G}\Omega_{i}^{\bm{s}}|\bm{s}_{i}^{t}-\bm{s}_{i}^{*}|+\Omega_{i}^{\bm{c}}|\bm{r}_{i}^{t}-\bm{r}_{i}^{*}|+\Omega_{i}^{d}|z_{i}^{t}-z_{i}^{*}|,$ (7)
where $G$ refers to the set of Gaussians involved in the optimization for the current frame; $\bm{s}_{i}^{t},\bm{r}_{i}^{t},z_{i}^{t}$ are scales, colors and depth to be optimized; $\bm{s}_{i}^{*},\bm{r}_{i}^{*},z_{i}^{*}$ are current parameter values. $\Omega$’s are importance weights for different parameters.
To calculate importance weights, for each Gaussian represented by Equation 1, we additionally add the following parameters:
$\Delta_{i}\mathrel{\mathop{:}}=\{N_{i}^{\mathrm{seen}},\Sigma_{i}^{\bm{s}},\Sigma_{i}^{\bm{c}},\Sigma_{i}^{\mathrm{d}}\}$ (8)
where $N_{i}^{\mathrm{seen}}$ records the “seen times” of each Gaussian; $\Sigma$ are the sum of gradients of $\mathcal{L}_{\mathrm{mapping}}$ with respect to parameters, representing the sensitivity to changes. After training each frame, we update the importance weights by averaging the gradients of mapping loss with respect to the parameters. Taking the scale s for example:
$\Omega_{i}^{\bm{s}}=\frac{\Sigma_{i}^{\bm{s}}}{N_{i}^{\mathrm{seen}}}=\frac{1}{N_{i}^{\mathrm{seen}}}\sum_{i}^{N_{i}^{\mathrm{seen}}}(\frac{\partial\mathcal{L}_{\mathrm{mapping}}}{\partial\bm{s}})$ (9)
The final loss function used in the mapping is
$\mathcal{L}_{\mathrm{{mapping}}}^{\prime}=\mathcal{L}_{\mathrm{mapping}}+\mathcal{L}_{\mathrm{reg}}$ (10)
IV-B Tracking
We conduct camera tracking in a render-and-compare way similar to that of iNeRF [3] for every input frame. The camera pose is parameterized by $T=(t,q)$, including the 3D translation $t$ and the rotation quaternion $q$. When tracking the new frame, we initialize the camera pose with the assumption of constant camera velocity. Since the color image can be affected by light, we decouple lightness from the image during tracking, similar to [22]. Specifically, let $\rho$ denote the operation of converting the image from the RGB space to the LAB space, and we discard the light channel (L channel) when comparing the two images.
$\mathcal{L}_{\mathrm{tracking}}=\lambda_{\mathrm{color}}|\rho(\hat{C})-\rho(C)|+\lambda_{\mathrm{depth}}|\hat{D}-D|$ (11)
V Experiments
In this section, we demonstrate through quantitative and qualitative experiments that our method achieves better reconstruction results than the current state-of-the-art neural implicit SLAM methods and concurrent work using Gaussian splatting. Furthermore, the ablation studies of section V-D show that our Gaussian densification strategy and regularization terms can effectively improve reconstruction quality.
V-A Experiment Setup
Evaluation Metrics: We use ATE (Absolute Trajectory Error [cm]) to measure the tracking performance, which aligns the estimated trajectory and the ground truth trajectory and then computes the root mean square error. To assess the quality of color image rendering, we employ peak signal-to-noise ratio (PSNR [dB]), structural similarity index (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS). PSNR evaluates pixel-wise RGB discrepancies, SSIM gauges structural resemblance, while LPIPS quantifies perceptual image patch similarity through neural network-based analysis. Since one can construct the mesh via TSDF-Fusion [23] with rendered depth along the estimated trajectory, we use metric Depth L1 [cm] to evaluate the geometric reconstruction as in [8]. All metrics are calculated by rendering full-resolution images along the estimated trajectory.
Datasets: We conduct our experiments on two datasets: Replica [24] (a synthetic dataset) and TUM-RGBD [25] (a real-world dataset); the synthetic dataset provides noiseless RGB and depth images while real-world datasets provide noisy inputs.
Baselines: We evaluate our results against both state-of-the-art SLAM methods using neural implicit representation —NICE-SLAM [4], ESLAM [5], Point-SLAM [8]—and 3D Gaussian Splatting—SplaTAM [11] and GS-SLAM [12].
V-B Reconstruction Performance
Quantitative Results.
TABLE I: The SLAM performance on the Replica [24] dataset. The best results are highlighted by first, second, and third. normal-$\uparrow$ means larger is better while normal-$\downarrow$ means smaller is better. Our method achieves the best results in most metrics.
Method Primitives Metric Room0 Room1 Room2 office0 office1 office2 office3 office4 Avg.
NICE-SLAM [4] Neural PSNR [dB] $\uparrow$ 22.12 22.47 24.52 29.07 30.34 19.66 22.23 24.94 24.42
+ SSIM $\uparrow$ 0.69 0.76 0.81 0.87 0.89 0.80 0.80 0.86 0.81
Voxels LPIPS $\downarrow$ 0.33 0.27 0.21 0.23 0.18 0.23 0.21 0.20 0.23
ATE RMSE [cm] $\downarrow$ 0.97 1.31 1.07 0.88 1.00 1.06 1.10 1.13 10.6
Depth L1 [cm] $\downarrow$ 1.81 1.44 2.04 1.39 1.76 8.33 4.99 2.01 2.97
[…]
+ SSIM $\uparrow$ 0.87 0.25 0.93 0.93 0.91 0.94 0.95 0.95 0.92
Feature Plane LPIPS $\downarrow$ 0.32 0.30 0.25 0.21 0.25 0.24 0.19 0.21 0.25
ATE RMSE [cm] $\downarrow$ 0.71 0.70 0.52 0.57 0.55 0.58 0.72 0.63 0.63
Depth L1 [cm] $\downarrow$ 0.97 1.07 1.28 0.86 1.26 1.71 1.43 1.06 1.18
[…]
+ SSIM $\uparrow$ 0.97 0.98 0.98 0.98 0.99 0.96 0.96 0.98 0.97
Point Cloud LPIPS $\downarrow$ 0.11 0.12 0.11 0.10 0.12 0.16 0.13 0.14 0.12
ATE RMSE [cm] $\downarrow$ 0.61 0.41 0.37 0.38 0.48 0.54 0.69 0.72 0.52
Depth L1 [cm] $\downarrow$ 0.53 0.22 0.46 0.30 0.57 0.49 0.51 0.46 0.44
GS-SLAM [12] PSNR [dB] $\uparrow$ 31.56 32.86 32.59 38.70 41.17 32.36 32.03 32.92 34.27
Parameterized SSIM $\uparrow$ 0.96 0.97 0.97 0.98 0.99 0.97 0.97 0.96 0.97
Gaussians LPIPS $\downarrow$ 0.09 0.07 0.09 0.05 0.03 0.09 0.11 0.11 0.08
ATE RMSE [cm] $\downarrow$ 0.48 0.53 0.33 0.52 0.41 0.59 0.46 0.70 0.50
Depth L1 [cm] 1.31 0.82 1.26 0.81 0.96 1.41 1.53 1.08 1.16
[…]
Parameterized SSIM $\uparrow$ 0.98 0.97 0.98 0.98 0.98 0.97 0.95 0.95 0.97
Gaussians LPIPS $\downarrow$ 0.07 0.10 0.08 0.09 0.09 0.10 0.12 0.15 0.10
ATE RMSE [cm] $\downarrow$ 0.31 0.40 0.29 0.47 0.27 0.29 0.32 0.55 0.36
Depth L1 [cm]
[…]
Parameterized SSIM $\uparrow$ 0.98 0.98 0.99 0.99 0.99 0.98 0.97 0.98 0.98
Gaussians LPIPS $\downarrow$ 0.05 0.05 0.04 0.03 0.03 0.07 0.08 0.08 0.05
ATE RMSE [cm] $\downarrow$ 0.19 0.34 0.16 0.21 0.26 0.23 0.21 0.38 0.25
Depth L1 [cm] $\downarrow$ 0.39 0.34 0.33 0.29 0.26 0.67 0.93 0.97 0.52
Compared to neural RGBD SLAM and more recent Gaussian Splatting SLAM, our method shows better rendering performance across both dataset, on real-world and synthetic data, as shown in table I and table II (it should be noted that since SplaTAM and GS-SLAM did not report rendering results on this dataset in their paper, we did not include them in the table).
Our method achieves the best results in color image rendering, but the second results in depth rendering in the Replica dataset. To explain why our method does not consistently outperform Point-SLAM on depth rendering, one must note that, in neural RGBD SLAM, appearance and geometry are often represented and optimized separately. In Point-SLAM, even though the same point cloud is used as primitives, each point is attached with a separate color feature descriptor and a geometric feature descriptor; also, two distinct neural networks are responsible for decoding color and depth. Hence, when optimizing the map parameters, color rendering and depth rendering do not affect each other. However, since our method uses Gaussian Splatting, the same Gaussian parameters $\{\mathbf{\mu},\mathbf{s},\mathbf{r},o\}$ are optimized for both color and depth rendering, which can impact the depth estimation.
Qualitative Results. Rendered images are shown in Figure 3. From the rendered results, Point-SLAM cannot recover fine details and SplaTAM generates floaters in the rendered image, which can be observed in the second and fourth row in fig. 3. On the other hand, our method generates clean images with high fidelity.
room 0
room2
Ground Truth Point-SLAM [8] SplaTAM [11] Ours
Figure 3: The rendering results on the Replica dataset. The second and forth rows are zoomed-in details of the colored squares. Compared to Point-SLAM [8], our method generates sharper results; compared to SplaTAM [11], our method doen not have floaters.
V-C Tracking Performance
On the Replica dataset, thanks to the improved reconstruction capabilities of our method shown in the previous section, we achieve the best tracking performance (see ATE RMSE in table I). However, on TUM-RGBD (table III), our method does not always perform the best. We hypothesise that, compared to methods using neural networks, our method is more sensitive to “noise”, such as motion blur and exposure changes present in real data.
TABLE II: The rendering performance on TUM-RGBD [25] dataset. The best results are highlighted by first, second, and third.
Method Metrics fr1/desk fr1/desk2 fr1/room fr3/off.
NICE-SLAM [4] PSNR [dB] $\uparrow$ 13.83 12.00 11.39 12.89
SSIM $\uparrow$ 0.57 0.51 0.37 0.55
LPIPS $\downarrow$ 0.48 0.52 0.63 0.50
[…]
SSIM $\uparrow$ 0.67 0.63 0.93 0.46
LPIPS 0.36 0.42 0.19 0.65
[…]
SSIM $\uparrow$ 0.63 0.59 0.65 0.75
LPIPS $\downarrow$ 0.54 0.57 0.55 0.45
GS-SLAM
[…]
SSIM $\uparrow$ 0.91 0.85 0.72 0.89
LPIPS $\downarrow$ 0.15 0.22 0.32 0.16
TABLE III: The tracking performance on TUM-RGBD [25] dataset. The best results are highlighted by first, second, and third. The evaluation metric for the trajectory is the ATE [cm].
Method fr1/desk fr1/desk2 fr1/room fr3/off.
NICE-SLAM [4] 4.26 4.99 34.49 3.87
ESLAM [5] 2.47 3.69 29.73 2.42
Point-SLAM [8] 4.34 4.54 30.92 3.48
GS-SLAM [12] 3.30 - - 6.60
SplaTAM [11] 3.35 6.54 11.13 5.16
Ours 3.38 7.20 22.62 5.12
V-D Ablation Study
In this section, we conduct ablation studies to demonstrate the effectiveness of the regularization terms and densification. For the regularization, we consistently render the first frame of Room0 in the Replica dataset at every step during the mapping. The fig. 4 shows rendered results of frame0 after the SLAM processing 350 frames. In the absence of regularization, the generated rendered image often exhibits artifacts such as floaters at the edge. Conversely, our approach ensures a cleaner output, with reduced or eliminated artifacts.
Rendered result after processing frame0 Rendered frame0 after 350 frames: Without Regularization Rendered frame0 after 350 frames: With Regularization
Figure 4: The difference of without/with Regularization. The first column shows the rendered result after just mapping frame0. The second column illustrates the rendered image of frame0 after processing 350 frames without regularization, while the third column showcases the same after applying regularization. The second rows show the zoomed-in results in the green square, we can see that with regularization, the result can still maintain good quality, especially for edges.
Quantitatively, we show the overall result on the Room0 in Replica dataset in table IV.
TABLE IV: Quantitative results of ablation study. We conduct ablation studies on Room0 in Replica dataset. “Regularization” means whether to add regularization terms (eq. 7) during mapping; “Densification” means whether to densify regions based on color and depth rendering. The results show our proposed methods can improve reconstruction quality and tracking accuracy.
Regularization Densification PSNR [dB] $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ ATE [cm] $\downarrow$
25.96 0.89 0.17 0.89
25.58 0.85 0.21 10.02
35.74 0.98 0.05 0.34
With regularization terms, we can achieve better reconstruction and tracking accuracy. Additionly, we also test with and without color and depth rendering densification, and the result (see the second row in table IV) reveals that the color and depth rendering densification can help improve rendering quality.
V-E Runtime
TABLE V: Runtime results on Room0 in Replica dataset.
Method Tracking [ms] / Iteration Mapping [ms] / Iteration Tracking [s] / Frame Mapping [s] / Frame
Point-SLAM [8] 31.50 19.05 0.63 7.62
SplaTAM [11] 75.25 91.66 3.01 5.50
Ours 45.00 65.57 1.80 4.59
Table V reports the runtime results on Room0 in Replica dataset running with a V100 GPU. For each iteration, Point-SLAM achieves the least time consumption because only a part of pixels are used in the tracking and mapping (in Point-SLAM, 1000 pixels are used in the mapping and 200 pixels are used in the tracking [^1] https://github.com/eriksandstroem/Point-SLAM). However, such a sparse sampling needs more iterations to converge, resulting in the longest time consumption for mapping; in addition, the sparse sampling decreases the rendering quality. SplaTAM and our method render full-resolution images, which require slightly more time at each iteration, but converge faster. It should be noted that our method is faster than SplaTAM due to implementation differences, since we render color and depth in one single rasterization process, as opposed to two rasterization processes in SplaTAM.
VI Conclusions
In this paper, we introduce a dense RGBD SLAM method based on the 3D Gaussian representation, which can render high-fidelity color and depth images. Our method uses rendering-based densification, which improves the quality of reobserved parts of the environment and allows the map to be extended. To alleviate the problem of “forgetting” (or overfitting) during mapping, we introduce a regularization loss in the optimization. Experiments show that our method consistently achieves higher-quality visual reconstruction than recent baselines, both those using neural representations and those using Gaussian representations. An interesting observation is that splatting-based methods, including ours, generally have lower tracking performance on real-world data (TUM-RGBD) compared to neural methods. However, our method still achieves the most high-fidelity visual reconstruction.
Despite the promising results we achieved in the experiments, our method has some limitations. As shown in table III, our method does not always achieve good tracking results in real-world datasets due to motion blur and varying exposure. As shown in fig. 5, though the rendered image is of higher quality since the mapping accumulates all previous frames, the blurry ground truth image makes tracking difficult. In addition, when evaluating the rendering quality, the fact that the ground truth image is blurry negatively affects the evaluation metric.
Ground truth image from TUM The rendering image PSNR: 17.16
Figure 5: The ground truth image and the rendering image of TUM dataset. Our method maps by accumulating all previous frames, leading to more visually pleasing image quality. However, the ground truth image is poor in this case due to motion blur and varying exposure, which brings difficulty to tracking. Also, it negatively affects the image similarity evaluation metric: PSNR is only 17.16 in this example.
To improve tracking accuracy, it is popular to use bundle adjustment to optimize camera poses and the map simultaneously. However, we experimentally found that simply adopting bundle adjustment as in BARF [26] in our experiments will only make the result worse. Motivated by [27], we plan to investigate loop closure detection and optimize the trajectory in a pose graph optimization (PGO) way in the future.
In addition to tracking accuracy, for future work, we will also investigate the following directions. First, how to improve efficiency to make it a “real-time” system. Second, motivated by [28, 29], we will investigate how to introduce high-level information, such as semantics, into the mapping.

### v2

@@
 To this end, we first propose a Gaussian densification strategy based on the rendering loss to map unobserved areas and refine reobserved areas.
-Second, we introduce extra regularization parameters to alleviate the “forgetting” problem in the continuous mapping problem, where parameters tend to overfit the latest frame and result in decreasing rendering quality for previous frames.
+Second, we introduce extra regularization parameters to alleviate the “forgetting” problem during contiunous mapping, where parameters tend to overfit the latest frame and result in decreasing rendering quality for previous frames.
 Both mapping and tracking are performed with Gaussian parameters by minimizing re-rendering loss in a differentiable way.
 Compared to recent neural and concurrently developed Gaussian splatting RGBD SLAM baselines, our method achieves state-of-the-art results on the synthetic dataset Replica and competitive results on the real-world dataset TUM.
+The code is released on https://github.com/ljjTYJR/HF-SLAM.
 Figure 1: Overview of the method.
@@
 Dense visual SLAM with RGBD inputs is essential for many downstream tasks in mobile robots, AR/VR and robot manipulation.
-Traditional SLAM systems have focused mainly on camera tracking [1] and geometric surface reconstruction [2].
+Traditional SLAM systems have focused mainly on camera tracking [mur2015orb] and geometric surface reconstruction [newcombe2011kinectfusion].
 However, surface appearance reconstruction, which contains rich information for scene understanding, is often lacking in traditional SLAM systems.
-Motivated by the success of NeRF [3] on novel view synthesis, recent RGBD SLAM methods use volumetric structures [4, 5, 6, 7] or points [8] as the map representation.
-Combined with neural networks as decoders, these methods can achieve both texture and geometry reconstruction.
-However, limited by the expensive sampling along the ray in the optimization of NeRF, the above mentioned methods cannot render full-resolution images with high quality (commonly only 0.5% pixels are used for mapping [5], leading to low quality image rendering; see fig.
-3 and table I).
-The recent work 3D Gaussian Splatting (3DGS) [9] provides a more efficient way for novel view synthesis.
+Motivated by the success of NeRF [yen2021inerf] on novel view synthesis, recent RGBD SLAM methods use volumetric structures [zhu2022nice, johari2023eslam, wang2023co, sucar2021imap] or points [sandstrom2023point] as the map representation.
+Combined with neural networks as decoders, these methods can achieve both appearance and geometry reconstruction.
+However, limited by the computationally expensive sampling along rays in the optimization of NeRF, the above mentioned methods cannot render full-resolution images with high quality (commonly only 0.5% pixels are used for mapping [johari2023eslam], leading to low quality image rendering; see Fig.
+3 and Tab.
+I).
+The recent work 3D Gaussian Splatting (3DGS) [kerbl20233dgaussian] provides a more efficient way for novel view synthesis.
 Instead of expensive sampling along the ray, 3DGS relies on rasterization for rendering, which accelerates the mapping process so as to handle full-resolution images.
-The original 3DGS needs prior camera poses, which are often estimated by Structure-from-Motion [10].
-In this work, we extend 3DGS in an SLAM framework, removing the need for prior camera poses.
+The original 3DGS needs prior camera poses, which are often estimated by Structure-from-Motion [schonberger2016structure].
+In this work, we extend 3DGS to the case of online tracking and mapping, removing the need for prior camera poses.
 Our method processes sequential RGBD frame inputs, simultaneously optimizing Gaussian parameters and estimating camera poses.
-Though there are some concurrent works combining 3DGS with SLAM [11, 12], we show our method achieves better or competitive results compared to recent baselines.
+Though there are some concurrent works combining 3DGS with SLAM [keetha2023splatam, yan2023gs], we show our method achieves better or competitive rendering and tracking results compared to recent baselines.
 Our first contribution is the introduction of a novel densification strategy based on rendering.
 We directly densify the map utilizing rendering loss, enabling us to effectively map unobserved areas and enhance the rendering quality of reobserved regions.
-Our second contribution is the regularized optimization during the mapping.
-The continuous mapping suffers from the “forgetting” problem, which means that Gaussian parameters tend to overfit the latest frame and lead to decrease of reconstruction quality for previous frames (cf.
-Figure 2).
-To alleviate the forgetting problem, we introduce extra parameters to supervise the learning process, and the experiment shows that our method can preserve the rendering quality of previously visited areas.
-Experiments show that our method achieves state-of-the-art reconstruction quality and tracking accuracy on Replica dataset and competitive results on the real-world dataset TUM-RGBD.
-In the ablation study, we show the effectiveness of our method.
-We also find that continuous mapping can reconstruct high-quality images even if there exists motion blur, which may inspire the 3DGS reconstruction with blurry inputs.
+Our second contribution is regularized optimization during mapping.
+Continuous mapping with 3DGS suffers from the “forgetting” problem, which means that Gaussian parameters tend to overfit to the latest frame and lead to decreased reconstruction quality for previous frames (see Fig.
+2).
+To alleviate the forgetting problem, we introduce extra parameters to supervise the learning process, and our experiments show that our method can preserve the rendering quality of previously visited areas.
+We demonstrate state-of-the-art reconstruction quality and tracking accuracy on Replica dataset and competitive results on the real-world dataset TUM-RGBD.
+We show the relative effectiveness of the two main contributions in the ablation study.
+We also find that continuous mapping can reconstruct high-quality images even if when faced with motion blur, which may inspire work on 3DGS reconstruction with blurry inputs.
 II Related Work
@@
 Conventional visual SLAM methods are divided into direct and indirect methods.
-Indirect methods (e.g., ORB-SLAM [1]) first detect points of interest and attach them by feature descriptors.
-The tracking is conducted by feature matching and minimizing the re-projection error across frames.
-Direct methods (e.g., DSO [13]) consume raw pixels and construct photometric loss for tracking.
-These conventional visual SLAM methods focus more on tracking instead of the scene reconstruction.
-Since NeRF [14] shows powerful scene reconstruction abilities, recent work applies neural representation in the SLAM to produce more complete and photorealistic scene reconstruction.
-For example, iMAP [7] uses a neural network to represent the scene; NICE-SLAM [4] and later work [5, 6] combine neural networks and voxel grids to improve reconstruction quality.
+Indirect methods (e.g., ORB-SLAM [mur2015orb]) first detect points of interest and attach feature descriptors.
+Tracking is conducted by feature matching and minimizing the re-projection error across frames.
+Direct methods (e.g., DSO [engel2017direct]) consume raw pixels and construct photometric loss for tracking.
+These conventional visual SLAM methods focus more on tracking instead of scene reconstruction.
+Since NeRF [mildenhall2021nerf] has shown powerful scene reconstruction abilities, recent work applies neural representation in SLAM to produce more complete and photorealistic scene reconstruction.
+For example, iMAP [sucar2021imap] uses a neural network to represent the scene; NICE-SLAM [zhu2022nice] and later work [johari2023eslam, wang2023co] combine neural networks and voxel grids to improve reconstruction quality.
 Thanks to the generalization capabilities of neural networks, these neural representations have more powerful abilities to fill holes and construct more complete scenes compared to explicit representations.
 Although neural representations achieve good results in geometric scene reconstruction, their performance in recovering high-fidelity scenes is not satisfactory enough.
-In this work, we use parameterized Gaussians [9] as the map representation and achieve a better photorealistic reconstruction.
-Concurrent with our work, some SLAM systems are also based on Gaussian Splatting: MonoGS [15] expands the map by random sampling; Gaussian-SLAM [16] divides the scene into many submaps; Photo-SLAM [17] conducts tracking by minimizing the reprojection error; SplaTAM [11] and GS-SLAM [12] densify the map only considering the unobserved regions.
+In this work, we use parameterized Gaussians [kerbl20233dgaussian] as the map representation and achieve a more photorealistic reconstruction.
+Concurrent with our work, some SLAM systems are also based on Gaussian Splatting: MonoGS [matsuki2023gaussian] expands the map by random sampling; Gaussian-SLAM [yugay2023gaussian] divides the scene into many submaps; Photo-SLAM [huang2023photo] conducts tracking by minimizing the reprojection error; SplaTAM [keetha2023splatam] and GS-SLAM [yan2023gs] densify the map only considering the unobserved regions (more information can be found in [chen2024survey]).
 Our method expands the map by considering filling holes in unobserved areas and refining reobserved areas, which can effectively improve the reconstruction quality.
@@
 II-B Photo-realistic reconstruction
-NeRF [14] opens a convenient way for photo-realistic reconstruction from only 2D images.
+NeRF [mildenhall2021nerf] opens a convenient way for photo-realistic reconstruction from only 2D images.
 Based on neural networks and volumetric differentiable rendering, NeRF can render novel photorealistic views that were not observed in the inputs.
-Many follow-up works promote original NeRF by accelerating training speed [18], improving rendering quality [19], etc.
-Some works [20] also reveal that neural networks are not necessary for novel view synthesis.
-As opposed to Gaussian Splatting [9], however, all the above-mentioned methods adopt expensive ray-marching when calculating one pixel, resulting in low rendering speed (lower than 15 fps).
-Gaussian Splatting uses parameterized Gaussians (as introduced in section III) as rendering primitives.
+Many follow-up works improve on the original NeRF by accelerating training speed [muller2022instant], improving rendering quality [barron2023zip], etc.
+Some works [fridovich2022plenoxels] also reveal that neural networks are not necessary for novel view synthesis.
+As opposed to Gaussian splatting [kerbl20233dgaussian], however, all the above-mentioned methods adopt expensive ray-marching when calculating one pixel, resulting in low rendering speed (lower than 15 fps).
+Gaussian Splatting uses parameterized Gaussians (as introduced in Sec.
+III) as rendering primitives.
 Instead of expensive ray-marching, Gaussian Splatting renders the image by tile-based rasterization, which can achieve much faster rendering speed (around 100 fps).
-The original Gaussian Splatting needs known camera poses (commonly achieved by SfM [10]) and optimizes Gaussian parameters offline.
+The original Gaussian Splatting needs known camera poses (commonly achieved by SfM [schonberger2016structure]) and optimizes Gaussian parameters offline.
 In this work, we consider sequential RGBD frame inputs and optimize parameters online, which releases the requirement of known camera poses and brings the possibility of online applications such as exploration.
 III Preliminary: Gaussian Splatting Rendering
-We use 3D Gaussians as the map representation primitives for the color and depth image rendering. As described in [9], each Gaussian consists of parameters
+We use 3D Gaussians as the map representation primitives for the color and depth image rendering. As described in [kerbl20233dgaussian], each Gaussian consists of parameters
 $\mathbf{\Theta}_{i}\mathrel{\mathop{:}}=\{\bm{\mu}_{i},\bm{s}_{i},\bm{r}_{i},\bm{c}_{i},o_{i}\},$ (1)
-where Gaussian means $\bm{\mu}_{i}\in\mathbb{R}^{3}$, scales $\bm{s_{i}}\in\mathbb{R}^{3}$, rotation quaternion $\bm{r}_{i}\in\mathbb{R}^{4}$, RGB color $\bm{c}_{i}\in\mathbb{R}^{3}$, and opacity $o_{i}\in\mathbb{R}$.
+with Gaussian means $\bm{\mu}_{i}\in\mathbb{R}^{3}$, scales $\bm{s_{i}}\in\mathbb{R}^{3}$, rotation quaternion $\bm{r}_{i}\in\mathbb{R}^{4}$, RGB color $\bm{c}_{i}\in\mathbb{R}^{3}$, and opacity $o_{i}\in\mathbb{R}$.
 All of these parameters are optimizable during training.
@@
 To render the color $C$ and depth $D$ of the selected pixel $x$ on the image plane, we blend $N$ ordered (by depth) 2D Gaussians overlapping the pixel:
-$C=\sum_{i\in N}\bm{c}_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j})$ (2)
-$D=\sum_{i\in N}z_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j})$ (3)
+$C=\sum_{i\in N}\bm{c}_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}),$ (2)
+$D=\sum_{i\in N}z_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}),$ (3)
 with
@@
 Though Eq.
-2 is the same as the volume rendering used in the classical NeRF [14], rasterization-based rendering is much more efficient.
-Because rasterization allows us to project and render objects in the space directly instead of expensive random sampling along the ray used in a full volumetric representation.
+2 is the same as the volume rendering used in the classical NeRF [mildenhall2021nerf], rasterization-based rendering is much more efficient.
+Because rasterization allows us to project and render objects in the space directly instead of expensive sampling along the ray used in a full volumetric representation.
 IV Methods
@@
 Given RGBD frames and estimated camera poses, we update the map by comparing the rendered images and the ground truth to identify unobserved regions and areas requiring refinement.
-Regularization terms are incorporated into the optimization process to mitigate the issue of forgetting during mapping (section IV-A).
-We track the camera pose in the Gaussian map by minimizing color and depth re-rendering loss (section IV-B).
+Regularization terms are incorporated into the optimization process to mitigate the issue of forgetting during mapping (Sec.
+IV-A).
+We track the camera pose in the Gaussian map by minimizing color and depth re-rendering loss (Sec.
+IV-B).
 IV-A Mapping
 Gaussian Densification.
-Different from initializing the Gaussian parameters with all images in the original 3D Gaussian Splatting [9], we process sequential RGBD frame inputs.
-For the latest RGBD input frame with an estimated pose (section IV-B), we need to add new Gaussians to the map, either to complete unobserved regions or to improve the rendering quality of previously mapped regions.
+Differently from initializing the Gaussian parameters with all images as in the original 3D Gaussian Splatting [kerbl20233dgaussian], we process sequential RGBD frame inputs.
+For the latest RGBD input frame with an estimated pose (Sec.
+IV-B), we need to add new Gaussians to the map, either to complete unobserved regions or to improve the rendering quality of previously mapped regions.
 First, for the unobserved regions, we need to add new Gaussians to fill holes.
@@
 That is, when receiving a new frame, we render the opacity image at the estimated pose by
-$O=\sum_{i\in N}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j})$ (5)
-In the opacity image, if the rendered value is below some predefined opacity threshold $\tau_{opa}$ (i.e., $O(x,y)<\tau_{\mathrm{opa}}$), we need to densify the corresponding regions to fill holes.
+$O=\sum_{i\in N}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}).$ (5)
+In the opacity image, if the rendered value is below some predefined opacity threshold $\tau_{\mathrm{opa}}$ (i.e., $O(x,y)<\tau_{\mathrm{opa}}$), we need to densify the corresponding regions to fill holes.
 Second, due to e.g.
-occlusion and illumination changes, the rendered image is often view-dependent; in addition, the optimized Gaussian parameters may get trapped into local minima to fit certain frames.
-To account for such view-dependent errors, we propose a kind of rendered error guided densification.
+occlusion and illumination changes, the rendered image is often view-dependent.
+In addition, the optimized Gaussian parameters may get trapped into local minima to fit certain frames.
+To account for such view-dependent errors, we propose a densification step guided by rendering error.
 When giving the estimated pose, we render the color image $\hat{C}$ and the depth image $\hat{D}$ at this pose; the current reference color image and the depth image are $C$ and $D$.
 We add new Gaussians when the error between the rendered image and the input image is large, where the color error and depth error are $E_{\mathrm{color}}(x,y)=|\hat{C}(x,y)-C(x,y)|$ and $E_{\mathrm{depth}}(x,y)=|\hat{D}(x,y)-D(x,y)|$ respectively.
-We will densify those regions where $E_{\mathrm{color}}(x,y)>\tau_{\mathrm{color}}$ or $\frac{E_{\mathrm{depth}}(x,y)}{D(x,y)}>\tau_{\mathrm{depth}}$, where $\tau_{\mathrm{color}}$ and $\tau_{\mathrm{depth}}$ are predefined thresholds.
+We will densify those regions where $E_{\mathrm{color}}(x,y)>\tau_{\mathrm{color}}$ or $\frac{E_{\mathrm{depth}}(x,y)}{D(x,y)}>\tau_{\mathrm{depth}}$, where $\tau_{\mathrm{color}}$ and $\tau_{\mathrm{depth}}$ are predefined thresholds—we divide by $D(x,y)$ when comparing to $\tau_{\text{depth}}$ to compensate for varying scales.
 In summary, we densify the regions where $\big{(}O(x,y)<\tau_{\mathrm{opa}}\big{)}\ |\ \big{(}E_{\mathrm{color}}(x,y)>\tau_{\mathrm{color}}\big{)}\ |\ \big{(}\frac{E_{\mathrm{depth}}(x,y)}{D(x,y)}>\tau_{\mathrm{depth}}\big{)}$.
-With depth images, we project corresponding depths to the space directly.
-The experiment (table IV) shows our densification strategy can significantly improve the map quality.
-Continuous Mapping.
-We optimize Gaussian parameters by minimizing the loss $\ell_{1}$ between the rendered color and depth image and the input frames; in addition, following the original implementation in [9], we add an extra $\mathrm{SSIM}$ term.
+With depth images, we project corresponding depths directly.
+The experiment reported in Tab.
+IV shows that our densification strategy can significantly improve map quality.
+[…]
 $\mathcal{L}_{\mathrm{mapping}}=$ $\lambda_{\mathrm{color}}|\hat{C}-C|+\lambda_{\mathrm{depth}}|\hat{D}-D|$
 $+\lambda_{\mathrm{SSIM}}\text{SSIM}(\hat{C},C)$ (6)
-As a continuous learning problem, the incremental mapping in our work also suffers from the forgetting problem [21].
-That is, the same Gaussian might be “seen” in multiple camera poses, but the Gaussian parameters tend to be optimized (overfitted) to fit the latest frame, resulting in a performance drop for previous frames (as demonstrated in fig.
+As a continuous learning problem, the incremental mapping in our work also suffers from the forgetting problem [mccloskey1989catastrophic]: the same Gaussian might be “seen” in multiple camera poses, and the Gaussian parameters tend to be optimized (overfitted) to fit the latest frame, resulting in a performance drop for previous frames (see Fig.
 2).
@@
 However, these Gaussians tend to be optimized to overfit the latest frame camera1, resulting in drop of reconstruction quality for previous frames.
-To alleviate the overfitting problem, prior work [4, 5, 6] maintains a keyframe buffer.
-When optimizing the latest frame, some previous keyframes are selected from the buffer based on some criteria such as overlap ratio with the current frame.
-The selected keyframes will then be optimized together with the current frame.
+To alleviate the overfitting problem, prior work [zhu2022nice, johari2023eslam, wang2023co] maintains a keyframe buffer from which previous keyframes are selected (based on criteria such as overlap ratio with the current frame) and the selected keyframes are then optimized together with the current frame.
 In the original 3D Gaussian Splatting, all frames are queued and picked randomly for many iterations to optimize Gaussian parameters.
-We can, of course, pick many previous frames whenever we optimize the current frame; however, optimizing with multiple frames will lead to more time consumption, which is not suitable for online continual mapping.
-To prevent Gaussian parameters from forgetting, we add an extra regularization term $\mathcal{L}_{\mathrm{reg}}$ when mapping:
+We can, of course, pick many previous frames whenever we optimize the current frame; however, optimizing with multiple frames leads to longer processing times, which is not suitable for online continual mapping.
+Instead, to prevent Gaussian parameters from forgetting, we add an extra regularization term $\mathcal{L}_{\mathrm{reg}}$ when mapping:
 $\mathcal{L}_{\text{reg}}=\sum_{i\in G}\Omega_{i}^{\bm{s}}|\bm{s}_{i}^{t}-\bm{s}_{i}^{*}|+\Omega_{i}^{\bm{c}}|\bm{r}_{i}^{t}-\bm{r}_{i}^{*}|+\Omega_{i}^{d}|z_{i}^{t}-z_{i}^{*}|,$ (7)
-where $G$ refers to the set of Gaussians involved in the optimization for the current frame; $\bm{s}_{i}^{t},\bm{r}_{i}^{t},z_{i}^{t}$ are scales, colors and depth to be optimized; $\bm{s}_{i}^{*},\bm{r}_{i}^{*},z_{i}^{*}$ are current parameter values.
+where $G$ refers to the set of Gaussians involved in the optimization for the current frame; $\bm{s}_{i}^{t},\bm{r}_{i}^{t},z_{i}^{t}$ are scales, colors and depth to be optimized at the step $t$; $\bm{s}_{i}^{*},\bm{r}_{i}^{*},z_{i}^{*}$ are current parameter values, that is, values achieved after step $t-1$.
 $\Omega$’s are importance weights for different parameters.
-To calculate importance weights, for each Gaussian represented by Equation 1, we additionally add the following parameters:
+In this way, we only need to maintain a small keyframe buffer but achieve better reconstruction quality.
+To calculate importance weights, for each Gaussian represented by Eq. 1, we additionally add the following parameters:
 $\Delta_{i}\mathrel{\mathop{:}}=\{N_{i}^{\mathrm{seen}},\Sigma_{i}^{\bm{s}},\Sigma_{i}^{\bm{c}},\Sigma_{i}^{\mathrm{d}}\}$ (8)
@@
 After training each frame, we update the importance weights by averaging the gradients of mapping loss with respect to the parameters.
-Taking the scale s for example:
-$\Omega_{i}^{\bm{s}}=\frac{\Sigma_{i}^{\bm{s}}}{N_{i}^{\mathrm{seen}}}=\frac{1}{N_{i}^{\mathrm{seen}}}\sum_{i}^{N_{i}^{\mathrm{seen}}}(\frac{\partial\mathcal{L}_{\mathrm{mapping}}}{\partial\bm{s}})$ (9)
+Taking the $i$th scale parameter $\textbf{s}_{i}$ as an example, after training with $N_{i}^{\mathrm{seen}}$ frames, its importance weight $\Omega_{i}^{\bm{s}}$ is:
+$\Omega_{i}^{\bm{s}}=\frac{\Sigma_{i}^{\bm{s}}}{N_{i}^{\mathrm{seen}}}=\frac{1}{N_{i}^{\mathrm{seen}}}\sum_{j}^{N_{i}^{\mathrm{seen}}}(\frac{\partial\mathcal{L}_{\mathrm{mapping}}}{\partial\bm{s}_{i}})_{j}$ (9)
 The final loss function used in the mapping is
-$\mathcal{L}_{\mathrm{{mapping}}}^{\prime}=\mathcal{L}_{\mathrm{mapping}}+\mathcal{L}_{\mathrm{reg}}$ (10)
+$\mathcal{L}_{\mathrm{{mapping}}}^{\prime}=\mathcal{L}_{\mathrm{mapping}}+\mathcal{L}_{\mathrm{reg}}.$ (10)
 IV-B Tracking
-We conduct camera tracking in a render-and-compare way similar to that of iNeRF [3] for every input frame.
+We conduct camera tracking in a render-and-compare way similar to that of iNeRF [yen2021inerf] for every input frame.
 The camera pose is parameterized by $T=(t,q)$, including the 3D translation $t$ and the rotation quaternion $q$.
-When tracking the new frame, we initialize the camera pose with the assumption of constant camera velocity.
-Since the color image can be affected by light, we decouple lightness from the image during tracking, similar to [22].
+When tracking the new frame, we initialize the camera pose assuming constant camera velocity.
+Since the color image can be affected by light, we decouple lightness from the image during tracking, similar to [wang2020self6d].
 Specifically, let $\rho$ denote the operation of converting the image from the RGB space to the LAB space, and we discard the light channel (L channel) when comparing the two images.
@@
 In this section, we demonstrate through quantitative and qualitative experiments that our method achieves better reconstruction results than the current state-of-the-art neural implicit SLAM methods and concurrent work using Gaussian splatting.
-Furthermore, the ablation studies of section V-D show that our Gaussian densification strategy and regularization terms can effectively improve reconstruction quality.
+Furthermore, the ablation studies of Sec.
+V-D show that our Gaussian densification strategy and regularization terms can effectively improve reconstruction quality.
 V-A Experiment Setup
@@
 PSNR evaluates pixel-wise RGB discrepancies, SSIM gauges structural resemblance, while LPIPS quantifies perceptual image patch similarity through neural network-based analysis.
-Since one can construct the mesh via TSDF-Fusion [23] with rendered depth along the estimated trajectory, we use metric Depth L1 [cm] to evaluate the geometric reconstruction as in [8].
-All metrics are calculated by rendering full-resolution images along the estimated trajectory.
-Datasets: We conduct our experiments on two datasets: Replica [24] (a synthetic dataset) and TUM-RGBD [25] (a real-world dataset); the synthetic dataset provides noiseless RGB and depth images while real-world datasets provide noisy inputs.
-Baselines: We evaluate our results against both state-of-the-art SLAM methods using neural implicit representation —NICE-SLAM [4], ESLAM [5], Point-SLAM [8]—and 3D Gaussian Splatting—SplaTAM [11] and GS-SLAM [12].
+Since one can construct the mesh via TSDF-Fusion [curless1996volumetric] with rendered depth along the estimated trajectory, we use metric Depth L1 [cm] to evaluate the geometric reconstruction as in [sandstrom2023point].
+All metrics are calculated by rendering full-resolution images along the estimated trajectory after reading all frames.
+Datasets: We conduct our experiments on two datasets: Replica [straub2019replica] (a synthetic dataset) and TUM-RGBD [sturm2012benchmark] (a real-world dataset); the synthetic dataset provides noiseless RGB and depth images while real-world datasets provide noisy inputs.
+Baselines: We evaluate our results against both state-of-the-art SLAM methods using neural implicit representation —NICE-SLAM [zhu2022nice], ESLAM [johari2023eslam], Point-SLAM [sandstrom2023point]—and 3D Gaussian Splatting—SplaTAM [keetha2023splatam] and GS-SLAM [yan2023gs].
+Implementation Details: We test all the methods on one V100 GPU with 24 GB memory.
+The hyper-parameters used in our method are as follows: $\tau_{\text{color}}=0.6$, $\tau_{\text{depth}}=0.1$, $\lambda_{\text{color}}=0.5$, $\lambda_{\text{depth}}=1.0$, $\lambda_{\text{SSIM}}=0.2$.
 V-B Reconstruction Performance
 Quantitative Results.
-TABLE I: The SLAM performance on the Replica [24] dataset.
+TABLE I: The SLAM performance on the Replica [straub2019replica] dataset.
 The best results are highlighted by first, second, and third.
-normal-$\uparrow$ means larger is better while normal-$\downarrow$ means smaller is better.
+$\uparrow$ means larger is better while $\downarrow$ means smaller is better.
 Our method achieves the best results in most metrics.
 Method Primitives Metric Room0 Room1 Room2 office0 office1 office2 office3 office4 Avg.
-NICE-SLAM [4] Neural PSNR [dB] $\uparrow$ 22.12 22.47 24.52 29.07 30.34 19.66 22.23 24.94 24.42
+NICE-SLAM [zhu2022nice] Neural PSNR [dB] $\uparrow$ 22.12 22.47 24.52 29.07 30.34 19.66 22.23 24.94 24.42
 + SSIM $\uparrow$ 0.69 0.76 0.81 0.87 0.89 0.80 0.80 0.86 0.81
@@
 Depth L1 [cm] $\downarrow$ 0.53 0.22 0.46 0.30 0.57 0.49 0.51 0.46 0.44
-GS-SLAM [12] PSNR [dB] $\uparrow$ 31.56 32.86 32.59 38.70 41.17 32.36 32.03 32.92 34.27
+GS-SLAM [yan2023gs] PSNR [dB] $\uparrow$ 31.56 32.86 32.59 38.70 41.17 32.36 32.03 32.92 34.27
 Parameterized SSIM $\uparrow$ 0.96 0.97 0.97 0.98 0.99 0.97 0.97 0.96 0.97
@@
 Depth L1 [cm] $\downarrow$ 0.39 0.34 0.33 0.29 0.26 0.67 0.93 0.97 0.52
-Compared to neural RGBD SLAM and more recent Gaussian Splatting SLAM, our method shows better rendering performance across both dataset, on real-world and synthetic data, as shown in table I and table II (it should be noted that since SplaTAM and GS-SLAM did not report rendering results on this dataset in their paper, we did not include them in the table).
+Compared to neural RGBD SLAM and more recent Gaussian Splatting SLAM, our method shows better rendering performance across both datasets, on real-world and synthetic data, as shown in Tab.
+I and Tab.
+II.
+(It should be noted that since SplaTAM and GS-SLAM did not report rendering results on this dataset in their paper, we did not include them in the table.)
 Our method achieves the best results in color image rendering, but the second results in depth rendering in the Replica dataset.
 To explain why our method does not consistently outperform Point-SLAM on depth rendering, one must note that, in neural RGBD SLAM, appearance and geometry are often represented and optimized separately.
-In Point-SLAM, even though the same point cloud is used as primitives, each point is attached with a separate color feature descriptor and a geometric feature descriptor; also, two distinct neural networks are responsible for decoding color and depth.
+In Point-SLAM, even though the same point cloud is used as primitives, each point has a separate color feature descriptor and a geometric feature descriptor; also, two distinct neural networks are responsible for decoding color and depth.
 Hence, when optimizing the map parameters, color rendering and depth rendering do not affect each other.
@@
 room2
-Ground Truth Point-SLAM [8] SplaTAM [11] Ours
+Ground Truth Point-SLAM [sandstrom2023point] SplaTAM [keetha2023splatam] Ours
 Figure 3: The rendering results on the Replica dataset.
 The second and forth rows are zoomed-in details of the colored squares.
-Compared to Point-SLAM [8], our method generates sharper results; compared to SplaTAM [11], our method doen not have floaters.
+Compared to Point-SLAM [sandstrom2023point], our method generates sharper results; compared to SplaTAM [keetha2023splatam], our method doen not have floaters.
 V-C Tracking Performance
-On the Replica dataset, thanks to the improved reconstruction capabilities of our method shown in the previous section, we achieve the best tracking performance (see ATE RMSE in table I).
-However, on TUM-RGBD (table III), our method does not always perform the best.
+On the Replica dataset, thanks to the improved reconstruction capabilities of our method shown in the previous section, we achieve the best tracking performance (see ATE RMSE in Tab.
+I).
+However, on TUM-RGBD (Tab.
+III), our method does not always perform the best.
 We hypothesise that, compared to methods using neural networks, our method is more sensitive to “noise”, such as motion blur and exposure changes present in real data.
-TABLE II: The rendering performance on TUM-RGBD [25] dataset. The best results are highlighted by first, second, and third.
+TABLE II: The rendering performance on TUM-RGBD [sturm2012benchmark] dataset. The best results are highlighted by first, second, and third.
 Method Metrics fr1/desk fr1/desk2 fr1/room fr3/off.
-NICE-SLAM [4] PSNR [dB] $\uparrow$ 13.83 12.00 11.39 12.89
+NICE-SLAM [zhu2022nice] PSNR [dB] $\uparrow$ 13.83 12.00 11.39 12.89
 SSIM $\uparrow$ 0.57 0.51 0.37 0.55
@@
 LPIPS $\downarrow$ 0.15 0.22 0.32 0.16
-TABLE III: The tracking performance on TUM-RGBD [25] dataset. The best results are highlighted by first, second, and third. The evaluation metric for the trajectory is the ATE [cm].
+TABLE III: The tracking performance on TUM-RGBD [sturm2012benchmark] dataset. The best results are highlighted by first, second, and third. The evaluation metric for the trajectory is the ATE [cm].
 Method fr1/desk fr1/desk2 fr1/room fr3/off.
-NICE-SLAM [4] 4.26 4.99 34.49 3.87
-ESLAM [5] 2.47 3.69 29.73 2.42
-Point-SLAM [8] 4.34 4.54 30.92 3.48
-GS-SLAM [12] 3.30 - - 6.60
-SplaTAM [11] 3.35 6.54 11.13 5.16
+NICE-SLAM [zhu2022nice] 4.26 4.99 34.49 3.87
+ESLAM [johari2023eslam] 2.47 3.69 29.73 2.42
+Point-SLAM [sandstrom2023point] 4.34 4.54 30.92 3.48
+GS-SLAM [yan2023gs] 3.30 - - 6.60
+SplaTAM [keetha2023splatam] 3.35 6.54 11.13 5.16
 Ours 3.38 7.20 22.62 5.12
@@
 For the regularization, we consistently render the first frame of Room0 in the Replica dataset at every step during the mapping.
-The fig.
-4 shows rendered results of frame0 after the SLAM processing 350 frames.
-In the absence of regularization, the generated rendered image often exhibits artifacts such as floaters at the edge.
+Figure 4 shows rendered results of frame0 after processing 350 frames.
+In the absence of regularization, the generated image often exhibits artifacts such as floaters at the edge.
 Conversely, our approach ensures a cleaner output, with reduced or eliminated artifacts.
@@
 The second rows show the zoomed-in results in the green square, we can see that with regularization, the result can still maintain good quality, especially for edges.
-Quantitatively, we show the overall result on the Room0 in Replica dataset in table IV.
+We show the overall quantitative results on the Room0 in Replica dataset in Tab. IV.
 TABLE IV: Quantitative results of ablation study.
@@
 With regularization terms, we can achieve better reconstruction and tracking accuracy.
-Additionly, we also test with and without color and depth rendering densification, and the result (see the second row in table IV) reveals that the color and depth rendering densification can help improve rendering quality.
+Additionally, we test with and without color and depth rendering densification, and the result (see the second row in Tab.
+IV) reveals that the color and depth rendering densification can help improve rendering quality.
 V-E Runtime
@@
 Method Tracking [ms] / Iteration Mapping [ms] / Iteration Tracking [s] / Frame Mapping [s] / Frame
-Point-SLAM [8] 31.50 19.05 0.63 7.62
-SplaTAM [11] 75.25 91.66 3.01 5.50
+Point-SLAM [sandstrom2023point] 31.50 19.05 0.63 7.62
+SplaTAM [keetha2023splatam] 75.25 91.66 3.01 5.50
 Ours 45.00 65.57 1.80 4.59
 Table V reports the runtime results on Room0 in Replica dataset running with a V100 GPU.
-For each iteration, Point-SLAM achieves the least time consumption because only a part of pixels are used in the tracking and mapping (in Point-SLAM, 1000 pixels are used in the mapping and 200 pixels are used in the tracking [^1] https://github.com/eriksandstroem/Point-SLAM).
+For each iteration, Point-SLAM achieves the least time consumption because only a subset of pixels are used in the tracking and mapping (in Point-SLAM, 1000 pixels are used in the mapping and 200 pixels are used in the tracking [^1] https://github.com/eriksandstroem/Point-SLAM).
 However, such a sparse sampling needs more iterations to converge, resulting in the longest time consumption for mapping; in addition, the sparse sampling decreases the rendering quality.
-SplaTAM and our method render full-resolution images, which require slightly more time at each iteration, but converge faster.
+SplaTAM and our method render full-resolution images, which requires slightly more time at each iteration, but converges faster.
 It should be noted that our method is faster than SplaTAM due to implementation differences, since we render color and depth in one single rasterization process, as opposed to two rasterization processes in SplaTAM.
-VI Conclusions
+VI Summary and Future Work
 In this paper, we introduce a dense RGBD SLAM method based on the 3D Gaussian representation, which can render high-fidelity color and depth images.
@@
 Despite the promising results we achieved in the experiments, our method has some limitations.
-As shown in table III, our method does not always achieve good tracking results in real-world datasets due to motion blur and varying exposure.
+As shown in Tab.
+III, our method does not always achieve good tracking results in real-world datasets due to motion blur and varying exposure.
 As shown in Fig.
@@
 Also, it negatively affects the image similarity evaluation metric: PSNR is only 17.16 in this example.
-To improve tracking accuracy, it is popular to use bundle adjustment to optimize camera poses and the map simultaneously.
-However, we experimentally found that simply adopting bundle adjustment as in BARF [26] in our experiments will only make the result worse.
-Motivated by [27], we plan to investigate loop closure detection and optimize the trajectory in a pose graph optimization (PGO) way in the future.
-In addition to tracking accuracy, for future work, we will also investigate the following directions.
-First, how to improve efficiency to make it a “real-time” system.
-Second, motivated by [28, 29], we will investigate how to introduce high-level information, such as semantics, into the mapping.
+[…]
+[…]
+[…]